OpenStudy (amina_soneviseth):

WILL FAN AND MEDAL BEST ANSWER FAN ALL ANSWERS!!! :D ------------------------------- -3/5x = 18 ------------------------------- y/2 = 8.7 how to solve it? ------------------------------- -1/5b = -3.5 how to solve it? ------------------------------ 14 = 2a/9

OpenStudy (amina_soneviseth):

@ArifL

OpenStudy (amina_soneviseth):

@Vocaloid

OpenStudy (anonymous):

it the still the same question?

OpenStudy (anonymous):

would u willing to solve it?

OpenStudy (anonymous):

To solve the second one you should multiply both sides by 2 and you will get your answer

OpenStudy (amina_soneviseth):

no, i understand it but i don't understand these problems because they are not whole numbers. @ArifL

OpenStudy (amina_soneviseth):

okay thank you! :D @kai47

OpenStudy (amina_soneviseth):

@ArifL i can try to solve it if you could help me find out how to solve the first one

OpenStudy (anonymous):

to solve the third one you would multiply each side by the reciprocal of -1/5 which is -5/1

OpenStudy (anonymous):

it just the same for the first one, multiply both side so it can be whole number

OpenStudy (amina_soneviseth):

@ArifL can u show me how to work the problem out?

OpenStudy (anonymous):

let see, \[-\frac{ 3 }{ 5 }x = 18\] multiply both side with 5 \[-\frac{ 3 }{ 5 }x(5) = 18(5)\]